# symbol/.style= draw=none, every to/.append style= edge node=node [sloped, allow upside down, auto=false]#$1$

### v1

@@ 6.1. Semiabelian varieties
Let $\pi_{A}:G\to A$ denote the maximal abelian quotient of $G$ and set $T:=\ker\pi_{A}$. We first claim that, for an arbitrary $\Sigma$-orbit $\widetilde{\Sigma}$, every $\widetilde{\Sigma}$-special subvariety of $A$ is $\widetilde{\Sigma}_{A}$-point-special and every $\widetilde{\Sigma}$-special subvariety of $T$ is $\widetilde{\Sigma}_{T}$-point-special. Indeed, suppose that $\widetilde{\Sigma}={\Sigma}(\widetilde{x})$ for some $\widetilde{x}\in\widetilde{G}_{0}(K)$ and that $W\subset T$ is $\widetilde{\Sigma}$-special. Because of Proposition 3.3, we can and will assume without loss of generality that $\widetilde{x}$ does not belong to any proper special subvariety of $\widetilde{G}_{0}$. By Lemma 5.4, there exist distinguished morphisms $\phi:Y\to\widetilde{G}_{0}$ and $\psi:Y\to T$ such that $W$ is an irreducible component of $\psi(\phi^{-1}(\widetilde{x}))$. As $\widetilde{x}$ does not belong to any proper special subvariety of $\widetilde{G}_{0}$, the morphism $\phi$ is surjective. We can and will assume that $\phi$ is a homomorphism of algebraic groups. Let $\pi^{\prime}:Y\to T_{Y}$ denote the maximal torus quotient of $Y$, which exists because of the Rosenlicht decomposition of $Y$ (see Proposition 3.1 in [5]), and let $\widetilde{\pi}:\widetilde{G}_{0}\to\widetilde{T}_{0}$ denote the maximal torus quotient of $\widetilde{G}_{0}$. By maximality, there exist a distinguished morphism $\chi:T_{Y}\to T$ and a homomorphism of algebraic groups $\omega:T_{Y}\to\widetilde{T}_{0}$ such that $\psi=\chi\circ\pi^{\prime}$ and $\widetilde{\pi}\circ\phi=\omega\circ\pi^{\prime}$. Furthermore, the surjectivity of $\phi$ implies that $Y/(\ker\phi+\ker\pi^{\prime})$ is a torus quotient of $\widetilde{G}_{0}$ and the maximality of $\widetilde{T}_{0}$ tells us that $Y/(\ker\phi+\ker\pi^{\prime})$ is a quotient of $\widetilde{T}_{0}$. Thus, $\ker(\widetilde{\pi}\circ\phi)\subset\ker\phi+\ker\pi^{\prime}$. The other inclusion is trivial and we have $\ker\phi+\ker\pi^{\prime}=\ker(\widetilde{\pi}\circ\phi)=\ker(\omega\circ\pi^{\prime})$. We deduce that
@@
Since $B$ is an abelian subvariety of $A_{0}\times A^{n}$, it follows that $(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})\subset B$. But $(y,\phi_{1}(y),\ldots,\phi_{n}(y))\in(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})(K)$ and so $B=(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})$. Combining this with the fact that $A_{0}$ and $A$ are the maximal abelian quotients of $G_{0}$ and $G$ respectively as well as with arguments from Subsection 4.1, we get
@@
6.2. Connected Shimura varieties
In the following Proposition 6.4 and its proof, we will use connected Shimura data and varieties and Shimura morphisms, submersions, immersions, coverings, and embeddings as well as quotient Shimura morphisms, quotient Shimura data, and Shimura subdata as defined in Definitions 3.2 and 3.4 in [3]. Given a connected mixed Shimura datum $(P,X^{+})$, we will also use the notation $P(\mathbb{Q})_{+}$ as introduced in Section 3 in [3].
@@ 6.2. Connected Shimura varieties
Let $S$ be a connected Shimura variety over $K$ and let $W\subset S$ be a weakly special subvariety. Then there exist connected Shimura varieties $S^{\prime}$, $S_{1}$, and $S_{2}$ as well as a Shimura embedding $i:S^{\prime}\to S$, a Shimura covering $p:S^{\prime}\to S_{1}\times S_{2}$, and a point $w\in S_{2}(K)$ such that $W$ is an irreducible component of $i(p^{-1}(S_{1}\times\{w\}))$.
@@
Suppose that $S=\mathcal{F}_{\mathrm{pSv}}(P,X^{+},\Gamma)$ where the functor $\mathcal{F}_{\mathrm{pSv}}$ is defined as in Section 3 in [3]. By Proposition 5.4 in [13], there exist connected Shimura varieties $S^{\prime}$ and $\widetilde{S}$, a Shimura embedding $i:S^{\prime}\to S$, a Shimura quotient morphism $\widetilde{p}:S^{\prime}\to\widetilde{S}$, and a point $\widetilde{w}\in\widetilde{S}(K)$ such that $W$ is an irreducible component of $i(\widetilde{p}^{-1}(\widetilde{w}))$. Proposition 5.4 in [13] holds over $\mathbb{C}$, but we can always find an algebraically closed subfield $L$ of $K$ over which $W$ is defined and such that $L$ embeds into $\mathbb{C}$. Proposition 5.4 in [13] then yields $\widetilde{w}\in\widetilde{S}(\mathbb{C})$, but since $i$ is finite by Proposition 3.6 in [3], we must have that $\widetilde{w}\in\widetilde{S}(L)$. Since $i$ is a finite morphism and since $\widetilde{p}^{-1}(\widetilde{w})$ is equidimensional by (A4), the above conclusion is equivalent to the fact ()($\ast$)( ∗ ) that $\widetilde{p}^{-1}(\widetilde{w})$ and $i^{-1}(W)$ share a common irreducible component. Suppose that $S^{\prime}=\mathcal{F}_{\mathrm{pSv}}(Q,Y^{+},\Delta)$. We can assume without loss of generality that $\widetilde{S}=\mathcal{F}_{\mathrm{pSv}}((Q,Y^{+})/N,\widetilde{\Delta})$ for some normal algebraic subgroup $N$ of $Q$ and some congruence subgroup $\widetilde{\Delta}\subset(Q/N)(\mathbb{Q})_{+}$ (see Remark 3.5 in [3]). Let $Z$ denote the center of $Q$, then we can form the quotient Shimura datum $(Q,Y^{+})/Z=:(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})$ as well as $(Q,Y^{+})/ZN$ (see again Remark 3.5 in [3]). Since $Z$ is of multiplicative type by [19, Corollary 17.62(a)], the two quotient Shimura morphisms $(Q,Y^{+})\to(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})$ and $(Q,Y^{+})/N\to(Q,Y^{+})/ZN$ are Shimura immersions and therefore also Shimura coverings. We now apply Lemma 2.2 in [21] to find a congruence subgroup $\Delta^{\prime}$ of $(Q/(ZN))(\mathbb{Q})$ that contains the image of $\widetilde{\Delta}$ under the map $(Q/N)(\mathbb{Q})\to(Q/(ZN))(\mathbb{Q})$. By Proposition 2.2 in [27] we may assume that $\Delta^{\prime}$ is contained in $(Q/(ZN))(\mathbb{Q})_{+}$. We replace $N$ by $ZN$, $\widetilde{\Delta}$ by $\Delta^{\prime}$, and $\widetilde{w}$ by its image under the induced Shimura morphism: thanks to (A4), this does not affect ()($\ast$)( ∗ ) or any other of the properties stated above. Having done that, we have that $Z\subset N$. Furthermore, by Theorem 21.51 in [19], we have that $Q^{\mathrm{ad}}=Q_{1}\times Q_{2}$ for some algebraic subgroups $Q_{1},Q_{2}$ of $Q^{\mathrm{ad}}$ such that $N/Z=Q_{1}\times\{1\}$. For $\{i,j\}=\{1,2\}$, we define connected Shimura data $(Q_{i},Y_{i}^{+})$ by $(Q_{i},Y_{i}^{+}):=(Q,Y^{+})/Q_{j}$. Note that $(Q_{2},Y_{2}^{+})\simeq(Q,Y^{+})/N$ and we will from now on identify these two connected Shimura data via the canonical isomorphism. We obtain a Shimura covering $(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})\to(Q_{1},Y_{1}^{+})\times(Q_{2},Y_{2}^{+})$. Let $\mathrm{ad}:Q\to Q^{\mathrm{ad}}$ and $\psi:Q^{\mathrm{ad}}\to Q/N=Q_{2}$ denote the quotient homomorphisms. Set $\Delta_{2}=\widetilde{\Delta}$ regarded as a congruence subgroup of $Q_{2}(\mathbb{Q})_{+}$. Fix any congruence subgroup $\Delta_{1}\subset Q_{1}(\mathbb{Q})_{+}$ (which exists, see Proposition 2.2 in [27]). Then $\Delta_{1}\times\Delta_{2}$ is a congruence subgroup of $Q^{\mathrm{ad}}(\mathbb{Q})_{+}$. Note that replacing $\Delta$ by some other congruence subgroup contained in $(\psi\circ\mathrm{ad})^{-1}(\Delta_{2})$ does not affect ()($\ast$)( ∗ ) or any other of the properties stated above. We can therefore again apply Lemma 2.2 in [21] and assume without loss of generality that $\Delta\subset\operatorname{ad}^{-1}(\Delta_{1}\times\Delta_{2})$. We now set $S_{i}=\mathcal{F}_{\mathrm{pSv}}(Q_{i},Y_{i}^{+},\Delta_{i})$ ($i=1,2$) and $p$ equal to the image under $\mathcal{F}_{\mathrm{pSv}}$ of the induced Shimura morphism $(Q,Y^{+},\Delta)\to(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}},\Delta_{1}\times\Delta_{2})\to(Q_{1},Y_{1}^{+},\Delta_{1})\times(Q_{2},Y_{2}^{+},\Delta_{2})$. We also let $w$ denote the point in $S_{2}(K)$ that is identified with $\widetilde{w}$ under the identification $((Q,Y^{+})/N,\widetilde{\Delta})=(Q_{2},Y_{2}^{+},\Delta_{2})$. The proposition now follows since $p^{-1}(S_{1}\times\{w\})=\widetilde{p}^{-1}(\widetilde{w})$. ∎
@@
In the distinguished category of connected (pure) Shimura varieties over $K$, the following holds. Let $\Sigma$ be the $\Sigma$-orbit of $x\in S_{0}(K)$. Then, a $\Sigma$-special subvariety of a distinguished variety $S$ contains a $\Sigma$-special point and is therefore $\Sigma_{S}$-point-special.
@@
Let $W$ be a $\Sigma$-special subvariety of $S$. By Proposition 6.4, there exist connected Shimura varieties $S^{\prime}$, $S_{1}$, and $S_{2}$ as well as a Shimura embedding $i:S^{\prime}\to S$, a Shimura covering $p:S^{\prime}\to S_{1}\times S_{2}$, and a point $w\in S_{2}(K)$ such that $W$ is an irreducible component of $i(p^{-1}(S_{1}\times\{w\}))$. It follows as in the proof of Proposition 6.4 that $i^{-1}(W)$ and $p^{-1}(S_{1}\times\{w\})$ share an irreducible component. By (A4), all irreducible components of $p^{-1}(S_{1}\times\{w\})$ have the same dimension, which must be equal to $\dim W$ as well as to $\dim S_{1}$ since $i$ and $p$ have finite fibers. By Proposition 5.11(3), this component is $\Sigma$-special in $S^{\prime}$, which implies by Proposition 5.11(1) that $\{w\}$ is $\Sigma$-special in $S_{2}$. By Lemma 5.6, we have that $w\in\Sigma_{S_{2}}$. By Lemma 13.3 in [18], whose proof can be modified to work also with our definition of a connected Shimura datum that is slightly different from the one in [18], there exists a special point $s_{1}\in S_{1}(K)$. By Propositions 3.4 and 3.5, we have that $(s_{1},w)\in\Sigma_{S_{1}\times S_{2}}$. So $S_{1}\times\{w\}$ is $\Sigma_{S_{1}\times S_{2}}$-point-special. By Proposition 5.11(3), any irreducible component of $p^{-1}(S_{1}\times\{w\})$ is $\Sigma_{S^{\prime}}$-point-special since $p$ is a Shimura covering, so has finite fibers and all irreducible components of $p^{-1}(S_{1}\times\{w\})$ are of dimension $\dim S_{1}$. By Proposition 5.11(1) and the above, $W$ is therefore $\Sigma_{S}$-point-special and we are done. ∎

### v2

@@ 6.1. Semiabelian varieties
Let $\pi_{A}:G\to A$ denote the maximal abelian quotient of $G$ and set $T:=\ker\pi_{A}$.
We first claim that, for an arbitrary $\Sigma$-orbit $\widetilde{\Sigma}$, every $\widetilde{\Sigma}$-special subvariety of $A$ is $\widetilde{\Sigma}_{A}$-point-special and every $\widetilde{\Sigma}$-special subvariety of $T$ is $\widetilde{\Sigma}_{T}$-point-special. Indeed, suppose that $\widetilde{\Sigma}={\Sigma}(\widetilde{x})$ for some $\widetilde{x}\in\widetilde{G}_{0}(K)$ and that $W\subset T$ is $\widetilde{\Sigma}$-special. Because of Proposition 3.3, we can and will assume without loss of generality that $\widetilde{x}$ does not belong to any proper special subvariety of $\widetilde{G}_{0}$. By Lemma 5.4, there exist distinguished morphisms $\phi:Y\to\widetilde{G}_{0}$ and $\psi:Y\to T$ such that $W$ is an irreducible component of $\psi(\phi^{-1}(\widetilde{x}))$. As $\widetilde{x}$ does not belong to any proper special subvariety of $\widetilde{G}_{0}$, the morphism $\phi$ is surjective. We can and will assume that $\phi$ is a homomorphism of algebraic groups. Let $\pi^{\prime}:Y\to T_{Y}$ denote the maximal torus quotient of $Y$, which exists because of the Rosenlicht decomposition of $Y$ (see Proposition 3.1 in [7]), and let $\widetilde{\pi}:\widetilde{G}_{0}\to\widetilde{T}_{0}$ denote the maximal torus quotient of $\widetilde{G}_{0}$. By maximality, there exist a distinguished morphism $\chi:T_{Y}\to T$ and a homomorphism of algebraic groups $\omega:T_{Y}\to\widetilde{T}_{0}$ such that $\psi=\chi\circ\pi^{\prime}$ and $\widetilde{\pi}\circ\phi=\omega\circ\pi^{\prime}$. Furthermore, the surjectivity of $\phi$ implies that $Y/(\ker\phi+\ker\pi^{\prime})$ is a torus quotient of $\widetilde{G}_{0}$ and the maximality of $\widetilde{T}_{0}$ tells us that $Y/(\ker\phi+\ker\pi^{\prime})$ is a quotient of $\widetilde{T}_{0}$. Thus, $\ker(\widetilde{\pi}\circ\phi)\subset\ker\phi+\ker\pi^{\prime}$. The other inclusion is trivial and we have $\ker\phi+\ker\pi^{\prime}=\ker(\widetilde{\pi}\circ\phi)=\ker(\omega\circ\pi^{\prime})$. We deduce that
@@
It now follows as in the proof of Proposition 6.1 that $W$ is $\widetilde{\Sigma}_{T}$-point-special. A completely analogous argument shows that every $\widetilde{\Sigma}$-special subvariety of $A$ is $\widetilde{\Sigma}_{A}$-point-special.
Let now $\pi_{0}:G_{0}\to A_{0}$ be the maximal abelian quotient of $G_{0}$ and $y=\pi_{0}(x)$. Because of Proposition 3.3, we can and will assume without loss of generality that $x$ does not belong to any proper special subvariety of $G_{0}$. We have seen in Subsection 4.1 that there are generators $\psi_{1},\dots,\psi_{n}$ of $\operatorname{\mathrm{Hom}}(G_{0},A)$ such that $\Sigma_{A}$ is equal to $\langle\gamma_{1},\dots,\gamma_{n}\rangle^{\mathrm{div}}$ where $\gamma_{i}=\psi_{i}(x)\in A(K)$ for all $i=1,\dots,n$. As all elements of $\operatorname{\mathrm{Hom}}(G_{0},A)$ factor through $A_{0}$, there are $\phi_{1},\dots,\phi_{n}\in\operatorname{\mathrm{Hom}}(A_{0},A)$ such that $\psi_{i}=\phi_{i}\circ\pi_{0}$ for all $i=1,\dots,n$. Therefore, $\gamma_{i}=\phi_{i}(y)$ for all $i=1,\dots,n$.
For each $i=1,\ldots,n$, fix $\gamma^{\prime}_{i}\in G(K)$ such that $\pi_{A}(\gamma^{\prime}_{i})=\gamma_{i}$. We claim that the conclusion of the proposition holds with $\Sigma^{\prime}=\Sigma(x^{\prime})$ for $x^{\prime}=(x,\gamma^{\prime}_{1},\ldots,\gamma^{\prime}_{n})\in(G_{0}\times G^{n})(K)$. Note that every $\Sigma$-special subvariety is $\Sigma^{\prime}$-special by Lemma 5.4; in particular, we have that $\Sigma_{Z}\subset\Sigma^{\prime}_{Z}$ for every semiabelian variety $Z$. We have just proved our first claim.
@@
Since $B$ is an abelian subvariety of $A_{0}\times A^{n}$, it follows that $(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})\subset B$. But $(y,\phi_{1}(y),\ldots,\phi_{n}(y))\in(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})(K)$ and so $B=(\operatorname{id}_{A_{0}},\phi_{1},\ldots,\phi_{n})(A_{0})$.
Combining this with the fact that $A_{0}$ and $A$ are the maximal abelian quotients of $G_{0}$ and $G$ respectively as well as with arguments from Subsection 4.1, we get
@@
6.2. Connected pure Shimura varieties
In the following Proposition 6.4 and its proof, we will use connected pure Shimura data and varieties and Shimura morphisms, submersions, immersions, coverings, and embeddings as well as quotient Shimura morphisms, quotient Shimura data, and Shimura subdata as defined in Definitions 3.1, 3.2, and 3.4 in [4]. Given a connected pure Shimura datum $(P,X^{+})$, we will also use the notation $P(\mathbb{Q})_{+}$ as introduced in Section 3 in [4].
@@ 6.2. Connected pure Shimura varieties
Let $S$ be a connected pure Shimura variety over $K$ and let $W\subset S$ be a weakly special subvariety. Then there exist connected pure Shimura varieties $S^{\prime}$, $S_{1}$, and $S_{2}$ as well as a Shimura embedding $i:S^{\prime}\to S$, a Shimura covering $p:S^{\prime}\to S_{1}\times S_{2}$, and a point $w\in S_{2}(K)$ such that $W$ is an irreducible component of $i(p^{-1}(S_{1}\times\{w\}))$.
@@
Suppose that $S=\mathcal{F}_{\mathrm{pSv}}(P,X^{+},\Gamma)$ where the functor $\mathcal{F}_{\mathrm{pSv}}$ is defined as in Section 3 in [4]. By Proposition 5.4 in [16], there exist connected pure Shimura varieties $S^{\prime}$ and $\widetilde{S}$, a Shimura embedding $i:S^{\prime}\to S$, a Shimura quotient morphism $\widetilde{p}:S^{\prime}\to\widetilde{S}$, and a point $\widetilde{w}\in\widetilde{S}(K)$ such that $W$ is an irreducible component of $i(\widetilde{p}^{-1}(\widetilde{w}))$ (note that Shimura subdata and quotient Shimura data of connected pure Shimura data are again pure by Proposition 2.9 in [16] and Remark 3.7 in [4]). Proposition 5.4 in [16] holds over $\mathbb{C}$, but we can always find an algebraically closed subfield $L$ of $K$ over which $W$ is defined and such that $L$ embeds into $\mathbb{C}$. Proposition 5.4 in [16] then yields $\widetilde{w}\in\widetilde{S}(\mathbb{C})$, but since $i$ is finite by Proposition 3.6 in [4], we must have that $\widetilde{w}\in\widetilde{S}(L)$. Since $i$ is a finite morphism and since $\widetilde{p}^{-1}(\widetilde{w})$ is equidimensional by (A4), the above conclusion is equivalent to the fact ()($\ast$)( ∗ ) that $\widetilde{p}^{-1}(\widetilde{w})$ and $i^{-1}(W)$ share a common irreducible component that surjects onto $W$ under $i$.
Suppose that $S^{\prime}=\mathcal{F}_{\mathrm{pSv}}(Q,Y^{+},\Delta)$. We can assume without loss of generality that $\widetilde{S}=\mathcal{F}_{\mathrm{pSv}}((Q,Y^{+})/N,\widetilde{\Delta})$ for some normal algebraic subgroup $N$ of $Q$ and some congruence subgroup $\widetilde{\Delta}\subset(Q/N)(\mathbb{Q})_{+}$ (see Remark 3.5 in [4]).
Let $Z$ denote the center of $Q$, then we can form the quotient Shimura datum $(Q,Y^{+})/Z=:(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})$ as well as $(Q,Y^{+})/ZN$ (see again Remark 3.5 in [4]). Since $Z$ is of multiplicative type by [22, Corollary 17.62(a)], the two quotient Shimura morphisms $(Q,Y^{+})\to(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})$ and $(Q,Y^{+})/N\to(Q,Y^{+})/ZN$ are Shimura immersions and therefore also Shimura coverings.
We now apply Lemma 2.2 in [24] to find a congruence subgroup $\Delta^{\prime}$ of $(Q/(ZN))(\mathbb{Q})$ that contains the image of $\widetilde{\Delta}$ under the map $(Q/N)(\mathbb{Q})\to(Q/(ZN))(\mathbb{Q})$. By Proposition 2.2 in [29] we may assume that $\Delta^{\prime}$ is contained in $(Q/(ZN))(\mathbb{Q})_{+}$. We replace $N$ by $ZN$, $\widetilde{\Delta}$ by $\Delta^{\prime}$, and $\widetilde{w}$ by its image under the induced Shimura morphism: thanks to (A4) and the fact that $(Q,Y^{+})/N\to(Q,Y^{+})/ZN$ is a Shimura covering, this does not affect ()($\ast$)( ∗ ) or any other of the properties stated above. Having done that, we have that $Z\subset N$.
Furthermore, by Theorem 21.51 in [22], we have that $Q^{\mathrm{ad}}=Q_{1}\times Q_{2}$ for some algebraic subgroups $Q_{1},Q_{2}$ of $Q^{\mathrm{ad}}$ such that $N/Z=Q_{1}\times\{1\}$. For $\{i,j\}=\{1,2\}$, we define connected pure Shimura data $(Q_{i},Y_{i}^{+})$ by $(Q_{i},Y_{i}^{+}):=(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})/Q_{j}$. Note that $(Q_{2},Y_{2}^{+})\simeq(Q,Y^{+})/N$ and we will from now on identify these two connected pure Shimura data via the canonical isomorphism. We obtain a Shimura covering $(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}})\to(Q_{1},Y_{1}^{+})\times(Q_{2},Y_{2}^{+})$.
Let $\mathrm{ad}:Q\to Q^{\mathrm{ad}}$ and $\psi:Q^{\mathrm{ad}}\to Q/N=Q_{2}$ denote the quotient homomorphisms. Set $\Delta_{2}=\widetilde{\Delta}$ regarded as a congruence subgroup of $Q_{2}(\mathbb{Q})_{+}$. Fix any congruence subgroup $\Delta_{1}\subset Q_{1}(\mathbb{Q})_{+}$ (which exists, see Proposition 2.2 in [29]). Then $\Delta_{1}\times\Delta_{2}$ is a congruence subgroup of $Q^{\mathrm{ad}}(\mathbb{Q})_{+}$. Note that replacing $\Delta$ by some other congruence subgroup contained in $(\psi\circ\mathrm{ad})^{-1}(\Delta_{2})$ does not affect ()($\ast$)( ∗ ) or any other of the properties stated above. We can therefore again apply Lemma 2.2 in [24] and assume without loss of generality that $\Delta\subset\operatorname{ad}^{-1}(\Delta_{1}\times\Delta_{2})$.
We now set $S_{i}=\mathcal{F}_{\mathrm{pSv}}(Q_{i},Y_{i}^{+},\Delta_{i})$ ($i=1,2$) and $p$ equal to the image under $\mathcal{F}_{\mathrm{pSv}}$ of the induced Shimura morphism $(Q,Y^{+},\Delta)\to(Q^{\mathrm{ad}},Y^{+,\mathrm{ad}},\Delta_{1}\times\Delta_{2})\to(Q_{1},Y_{1}^{+},\Delta_{1})\times(Q_{2},Y_{2}^{+},\Delta_{2})$. We also let $w$ denote the point in $S_{2}(K)$ that is identified with $\widetilde{w}$ under the identification $((Q,Y^{+})/N,\widetilde{\Delta})=(Q_{2},Y_{2}^{+},\Delta_{2})$. The proposition now follows since $p^{-1}(S_{1}\times\{w\})=\widetilde{p}^{-1}(\widetilde{w})$. ∎
@@
In the distinguished category of connected pure Shimura varieties over $K$, the following holds. Let $\Sigma$ be the $\Sigma$-orbit of $x\in S_{0}(K)$. Then, a $\Sigma$-special subvariety of a connected pure Shimura variety $S$ contains a $\Sigma$-special point and is therefore $\Sigma_{S}$-point-special.
@@
Let $W$ be a $\Sigma$-special subvariety of $S$. By Proposition 6.4, there exist connected pure Shimura varieties $S^{\prime}$, $S_{1}$, and $S_{2}$ as well as a Shimura embedding $i:S^{\prime}\to S$, a Shimura covering $p:S^{\prime}\to S_{1}\times S_{2}$, and a point $w\in S_{2}(K)$ such that $W$ is an irreducible component of $i(p^{-1}(S_{1}\times\{w\}))$. By (A4), all irreducible components of $p^{-1}(S_{1}\times\{w\})$ have the same dimension, which must be equal to $\dim W$ as well as to $\dim S_{1}$ since $i$ and $p$ have finite fibers. It follows as in the proof of Proposition 6.4 that $i^{-1}(W)$ and $p^{-1}(S_{1}\times\{w\})$ share an irreducible component. By Proposition 5.11(3), this component is $\Sigma$-special in $S^{\prime}$, which implies by Proposition 5.11(1) that $\{w\}$ is $\Sigma$-special in $S_{2}$. By Lemma 5.6, we have that $w\in\Sigma_{S_{2}}$. By Lemma 13.3 in [21], whose proof can be modified to work also with our definition of a connected Shimura datum that is slightly different from the one in [21] (see Remark 4.4 in [4]), there exists a special point $s_{1}\in S_{1}(K)$. By Propositions 3.4 and 3.5, we have that $(s_{1},w)\in\Sigma_{S_{1}\times S_{2}}$. So $S_{1}\times\{w\}$ is $\Sigma_{S_{1}\times S_{2}}$-point-special. By Proposition 5.11(3), every irreducible component of $p^{-1}(S_{1}\times\{w\})$ is $\Sigma_{S^{\prime}}$-point-special since $p$ is a Shimura covering, so has finite fibers and all irreducible components of $p^{-1}(S_{1}\times\{w\})$ are of dimension $\dim S_{1}$. By Proposition 5.11(1) and the above, $W$ is therefore $\Sigma_{S}$-point-special and we are done. ∎
Our notion of a $\Sigma$-special subvariety of a connected pure Shimura variety $S$ (which has now turned out to coincide with the notion of “$\Sigma_{S}$-point-special”) is related to the sets $WS_{X}(M,X_{M},x)$ defined in [31], see Section 5 and Theorem 7.1 in [31] and Lemma 5.4.